# Fourier-Mukai Partners of Abelian Varieties and K3 surfaces in Positive and Mixed Characteristics

### v1

@@ 1. Introduction
One of the first attempts to understand the landscape of Fourier-Mukai equivalences in positive characteristics is due to Lieblich and Olsson ([LO15]), where a key technique was the lifting of Fourier-Mukai transforms (or their kernels) from characteristic $p$ to characteristic 00. Motivated by this, a natural question is whether there are interesting classes of examples where one can lift Fourier-Mukai equivalences from characteristic $p$ to characteristic 00. In [Del81a], starting with an ordinary abelian variety or an ordinary K3 surface (the precise assumptions are as in Theorem 1.1 below), Deligne exhibits the so-called “canonical coordinates” on its mixed-characteristic formal moduli space and uses them to constructs its (algebraizable) canonical lift over the ring of Witt vectors (see also [Del81b], [Kat81], [BT24]). The first contribution of this paper is that such canonical lifts provide examples where kernels of Fourier-Mukai equivalences can be always lifted, answering the question above:
Theorem 1.1 (Lifting Fourier-Mukai equivalences).
@@
Theorem 1.5 (Relative Fourier-Mukai partners of canonical lifts).
@@
Corollary 1.6 (Special case of [HLT20, Theorem 1.1]).
@@
Theorem 1.7 (Fourier-Mukai partners of abelian varieties).
@@
Assuming for a moment that $k$ is algebraically closed, if the answer to Question 1.10 is yes, it will imply that $\operatorname{FMP}(X_{\operatorname{can}}/W(k))$ is a countable set, as this is known for $\operatorname{FMP}(X/k)$ ([Sta, Tag 0G11], [AT09]). After completing the draft for this paper, the author found a direct proof of this countability result ([Kur24]) [^1] We are considering only smooth projective schemes over the ring of Witt vectors here, but the main result of [Kur24] is much more general and concerns smooth proper schemes over a noetherian base., so it provides some positive evidence toward the question.
@@ Acknowledgement
The author would like to thank his advisor Alexander Perry for continual support, helpful discussions and suggestions. The author thanks Sean Cotner, Hyunsuk Kim, Martin Olsson, Saket Shah and Gleb Terentiuk for helpful discussions and interest. The author also thanks Sean Cotner for suggesting a slight simplification in the proof of Proposition 3.3.
@@ 2. Relative Fourier-Mukai transform
Remark 2.4.
In this paper, we will only consider smooth projective $S$-schemes in the context of Fourier-Mukai partners.
@@ 3. Fourier-Mukai partners of abelian varieties
Proposition 3.1 (Combination of [Rou11, Théoréme 4.18] and [PS11, Corollary B]).
@@
The paper [PS11] works over the complex numbers, but this proposition as stated is still true in any characteristic by the same proof. (Note that in positive characteristics, the Hodge number $h^{1,0}(X)$ may differ from the dimension of $\operatorname{Pic}_{X/k}^{0}$.)
@@ 4.2. Grothendieck-Messing theory via Mukai structures
Grothendieck-Messing theory describes lifts of abelian varieties via the data of Hodge filtrations. In this subsection, specializing to the case of canonical lifts, we rewrite the theory in terms of the Mukai $F$-isocrystals. We assume throughout the subsection that $k$ is an algebraically closed field of characteristic $p>0$, and we will use the notations $W=W(k),K=\operatorname{Frac}W$. We briefly recall the notions of $F$-(iso)crystals.
@@
If $X$ is a smooth proper $k$-variety, for each $i$, the crystalline cohomology modulo torsion $H_{\rm cris}^{i}(X/W)/(\rm torsion)$ with its Frobenius action is an $F$-crystal over $k$, and $H^{i}_{\rm cris}(X/K)$ is an $F$-isocrystal over $k$, where the injectivity condition of $F$-(iso)crystals follows from Poincaré duality.
@@ 4.3. Proofs of the main results for abelian varieties
Let $k$ be an algebraically closed positive characteristic field. The Newton polygon of an abelian variety over $k$ is derived-invariant. In particular, ordinarity and supersingularity of abelian varieties are derived-invariant.
@@
Let $A,B$ be two derived-equivalent abelian varieties over $k$. We know that there exists an isomorphism of abelian varieties $A\times\widehat{A}\simeq B\times\widehat{B}$ (see Proposition 3.1). In particular, we have
@@
as $F$-crystals. Since an abelian variety is isogenous to its dual abelian variety, they have the same Newton polygons. Thus, the isomorphism of $F$-crystals above shows that $A$ and $B$ have the same Newton polygons.
The Hodge polygon of an abelian variety is determined by its dimension, so ordinarity and supersingularity are derived-invariant ([CL98, III Théorème 3.2]). ∎
Lemma 4.18.
Let $A$ be an ordinary abelian variety. If $B$ is another abelian variety which is either isogenous to $A$ or a sub-abelian variety of $A$, then $B$ is also ordinary.
Proof.
We recall that the Hodge polygon of an abelian variety is determined by its dimension. Since the Newton polygon is isogeny-invariant, the case of an isogeny follows from [CL98, III Théorème 3.2]. If $B$ is a sub-abelian variety of $A,$ we have a short exact sequence of abelian varieties
${e\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${B\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota}$${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${A/B\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${e}$
This splits up to isogeny as:
${A/B\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[m]}$$\scriptstyle{s}$${e\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${B\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota}$${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${A/B\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${e}$
We claim that $(\iota,s):B\times A/B\longrightarrow A$ is an isogeny. Since the source and the target have the same dimensions, it suffices to show that the map is surjective, or that it’s surjective on $k$-points. Given some $x\in A(k)$, there is some $y\in A/B(y)$ with $\pi(x)=[m](y)$. Then, $\pi(x-s(y))=\pi(x)-[m](y)=e_{A/B}$, so $x-s(y)\in B(k)$. We have
$(\iota,s)(x-s(y),y)=x,$
so we see that $B\times A/B\longrightarrow A$ is surjective. Now, the Newton polygon of $A$ is the same as the Newton polygon of $B\times A/B$, so $B$ must be ordinary in order for $A$ to be ordinary. ∎
The key ingredient for Theorem 1.1 in the abelian variety case is the classification of point objects on abelian varieties. We start with the version in characteristic 00:
Proposition 4.19 (Special case of one implication in [dJO22] for algebraically closed fields).
Let $L$ be an algebraically closed field of characteristic 0, and let $X$ be an abelian variety over $L$ of dimension $d$. Let $F\in\operatorname{D}^{\operatorname{b}}_{\operatorname{coh}}(X)$ be an object satisfying
@@
Then, there is a sub-abelian variety $i:Y\hookrightarrow X$, a rational point $x\in X(L)$, a simple semi-homogeneous vector bundle $V$ on $Y$, and an integer $r$ such that $F\simeq(t_{x})_{*}i_{*}V[r]$, where $t_{x}:X\longrightarrow X$ is the translation map by $x$.
The proof of this result appears to use the characteristic assumption in several places (cf. [dJO22, 3.9, 6.1, 6.2]), but as [dJO22, 1.3] suggests, the proof can be adopted to positive characteristics, except the step [dJO22, 6.1]. For our purposes, we will use a variant of their result which works in any characteristics. Let $L$ be an algebraically closed field of any characteristic, and let $X$ be an abelian variety over $L$ of dimension $d$. Let $s\mathcal{D}$ be the moduli stack of simple relatively perfect universally gluable complexes for the morphism $X\longrightarrow\operatorname{Spec}L$. This is an algebraic stack quasi-separated and locally of finite type over $L$ ([Lie06], [Sta, Tag 0DPX]), and it’s a ${\mathbb{G}_{m}}$-gerbe over the coarse space $sD$ which is an algebraic space quasi-separated and locally of finite type over $L$ ([LMB00, Corollaire 10.8]). Given a simple gluable complex $F\in\operatorname{D}^{\operatorname{b}}_{\operatorname{coh}}(X)$, it defines an $L$-point $[F]:\operatorname{Spec}L\longrightarrow s\mathcal{D}$, or an $L$-point $[F]:\operatorname{Spec}L\longrightarrow sD$. Similarly to [dJO22, 5.8], we consider the action map of $X\times_{L}\widehat{X}$ on $sD$ denoted as $\mu_{F}:X\times_{L}\widehat{X}\longrightarrow sD$ which acts on the $L$-points as $(x,L)\mapsto(t_{x}^{*}F)\otimes L$. We define the stabilizer $S_{F}$ by the cartesian square:
@@
By the following lemma, $S_{F}$ is in fact a closed subgroup variety of $X\times_{L}\widehat{X}$:
@@
With this preparation, essentially the same argument as [dJO22] shows:
@@
This vanishes for $i<0$ and has dimension 1 for $i=0$ (Note that these do not require the HKR isomorphisms to hold), so the first two conditions are satisfied. For the last condition, we are looking at the stabilizer of $E$ in $(A\times_{k}B)\times_{k}(\widehat{A}\times_{k}\widehat{B})$. [Orl03, Corollary 2.13] shows that a rational point $(a,b,\alpha,\beta)\in(A\times_{k}B\times_{k}\widehat{A}\times_{k}\widehat{B})(k)$ satisfies $f_{E}(a,\alpha)=(b,\beta)$ if and only if $(a,b,\alpha,-\beta)$ is in the stabilizer of $E$, where $f_{E}:A\times_{k}\widehat{A}\simeq B\times_{k}\widehat{B}$ is the isomorphism as in [Orl03, Theorem 2.10]. In other words, the reflection of $(S_{F})_{\operatorname{red}}$ on the $\widehat{B}$-component coincides with the graph $\Gamma_{f_{E}}$ of the map $f_{E}$. This has dimension $2g$, so we are done. ∎
@@
The proposition implies in particular that if $A$ and $B$ are derived-equivalent ordinary abelian varieties over $k$, then $A_{\operatorname{can},\overline{\eta}},B_{\operatorname{can},\overline{\eta}}$ are also derived-equivalent.
@@
When $\mathcal{X}=A_{\operatorname{can}}$ for an ordinary abelian variety $A$, the condition in the conjecture regarding the Hodge filtration is vacuous. To see this, by the compatibility of Chern classes with cycle classes ([Gro58, (16)]), it suffices to check that the crystalline Chern classes always land in the middle Hodge filtration under the identification $H_{\rm cris}^{*}(A/W)\simeq H_{\rm dR}^{*}(A_{\operatorname{can}}/W)$. By the proof of Proposition 4.16, the isomorphism identifies the Hodge filtration with the slope filtration, but the $i$-th crystalline Chern class always lands in the $i$-th filtered piece of the slope filtration ([LO15, §2.2]), so we have verified the condition. Lemma 4.22 yields the following special case of this conjecture (which can be applied to kernels of Fourier-Mukai equivalences as in the proof of Proposition 4.23):
@@
[…]

### v2

@@ Abstract.
footnotetext: Keywords: derived equivalence, abelian varieties, K3 surfaces, positive and mixed characteristics.footnotetext: MSC2020: 14F08, 14G17, 14G45, 14J28, 14K05.
@@ 1. Introduction
One of the first attempts to understand the landscape of Fourier-Mukai equivalences in positive characteristics is due to Lieblich and Olsson ([LO15]), where a key technique was the lifting of Fourier-Mukai transforms (or their kernels) from characteristic $p$ to characteristic 0. Motivated by this, a natural question is whether there are interesting classes of examples where one can lift Fourier-Mukai equivalences from characteristic $p$ to characteristic 0. In [Del81a], starting with an ordinary abelian variety or an ordinary K3 surface (the precise assumptions are as in Theorem 1.1 below), Deligne exhibits the so-called “canonical coordinates” on its mixed-characteristic formal moduli space and uses them to constructs its (algebraizable) canonical lift over the ring of Witt vectors (see also [Del81b], [Kat81], [BT24]). The first contribution of this paper is that such canonical lifts provide examples where kernels of Fourier-Mukai equivalences can be always lifted, answering the question above:
Theorem 1.1 (Lifting Fourier-Mukai equivalences; Propositions 4.23, 5.6).
@@
Theorem 1.5 (Relative Fourier-Mukai partners of canonical lifts; Propositions 4.30, 5.8).
@@
Corollary 1.6 (Special case of [HLT20, Theorem 1.1]; Proposition 6.1).
@@
Theorem 1.7 (Fourier-Mukai partners of abelian varieties; Proposition 3.3).
@@
Assuming for a moment that $k$ is algebraically closed, if the answer to Question 1.10 is yes, it will imply that $\operatorname{FMP}(X_{\operatorname{can}}/W(k))$ is a countable set, as this is known for $\operatorname{FMP}(X/k)$ ([Sta, Tag 0G11], [AT09]). After completing the draft for this paper, the author found a direct proof of this countability result ([Kur24])We are considering only smooth projective schemes over the ring of Witt vectors here, but the main result of [Kur24] is much more general and concerns smooth proper schemes over a noetherian base., so it provides some positive evidence toward the question.
@@ Acknowledgement
The author would like to thank his advisor Alexander Perry for continual support, helpful discussions and suggestions. The author thanks Sean Cotner, Hyunsuk Kim, Martin Olsson, Saket Shah and Gleb Terentiuk for helpful discussions and interest. The author also thanks Sean Cotner for suggesting a slight simplification in the proof of Proposition 3.3. The author thanks the reviewer for helpful comments.
@@ 2. Relative Fourier-Mukai transform
Remark 2.4.
While the above definition applies to any smooth proper schemes over $S$, we will only consider Fourier-Mukai partners that are smooth projective over $S$ in this paper.
@@ 3. Fourier-Mukai partners of abelian varieties
Proposition 3.1 (Adaptation of [Rou11, Théoréme 4.18] and [PS11, Corollary B] to all characteristics).
@@
The first statement of the proposition is due to [Rou11], and the rest of the proposition is the adaptation of [PS11, Corollary B]. The paper [PS11] works over the complex numbers and proves $h^{0}(X,\Omega^{1}_{X})=h^{0}(Y,\Omega^{1}_{Y})$ in Corollary B as a restatement of $\dim\operatorname{Pic}^{0}_{X/k}=\dim\operatorname{Pic}^{0}_{Y/k}$. However, their proof of $\dim\operatorname{Pic}^{0}_{X/k}=\dim\operatorname{Pic}^{0}_{Y/k}$ works in all characteristics, so our proposition as stated is still true in any characteristic. In positive characteristics, the Hodge number $h^{1,0}(X)$ can differ from the dimension of $\operatorname{Pic}_{X/k}^{0}$, and in fact Addington and Bragg constructed derived equivalent smooth projective threefolds $X,Y$ over $\overline{\mathbf{F}}_{3}$ with $h^{0}(X,\Omega^{1}_{X})\neq h^{0}(Y,\Omega^{1}_{Y})$ in [ABP23].
@@ 4.2. Grothendieck-Messing theory via Mukai structures
Grothendieck-Messing theory describes lifts of abelian varieties via the data of Hodge filtrations. In this subsection, specializing to the case of canonical lifts, we rewrite the theory in terms of the Mukai $F$-isocrystals. We assume throughout the subsection that $k$ is an algebraically closed field of characteristic $p>0$, and we will use the notations $W=W(k),K=\operatorname{Frac}W$. We briefly recall the notions of FF-(iso)crystals.
@@
If $X$ is a smooth proper $k$-variety, for each $i$, the crystalline cohomology modulo torsion $H_{\rm cris}^{i}(X/W)/(\rm torsion)$ with its Frobenius action is an $F$-crystal over $k$, and $H^{i}_{\rm cris}(X/K)$ is an $F$-isocrystal over $k$, where the injectivity condition of FF-(iso)crystals follows from Poincaré duality.
@@ 4.3. Proofs of the main results for abelian varieties
Let $A$ and $B$ be abelian varieties over an algebraically closed field of positive characteristic. In any of the following situations, if $A$ is ordinary (resp. supersingular), so is $B$.
  1. (1)
    $B$ is isogenous to $A$.
  2. (2)
    $B$ is derived equivalent to $A$.
  3. (3)
    $B$ is a sub-abelian variety of $A$.
@@
We recall that the Hodge polygon of an abelian variety is determined by its dimension. Since the Newton polygon is isogeny-invariant, the case of an isogeny follows from [CL98, III Théorème 3.2].
If $B$ is derived equivalent to $A$, there exists an isomorphism of abelian varieties $A\times\widehat{A}\simeq B\times\widehat{B}$ (see Proposition 3.1). In particular, we have
@@
as $F$-crystals. Since an abelian variety is isogenous to its dual abelian variety, they have the same Newton polygons. Thus, the isomorphism of $F$-crystals above shows that $A$ and $B$ have the same Newton polygons. This treats the second case.
Finally, if $B$ is a sub-abelian variety of $A,$ then $B\times(A/B)$ and $A$ are isogenous. Since the Newton polygon of $A$ is the same as the Newton polygon of $\times(A/B), BB$ must be ordinary (resp. supersingular) in order for $A$ to be ordinary (resp. supersingular). ∎
The key ingredient for Theorem 1.1 in the abelian variety case is the classification of point objects on abelian varieties. This classification is a variant of the following result in characteristic 0:
Proposition 4.18 ([dJO22, Theorem 1.1] over an algebraically closed field).
Let $L$ be an algebraically closed field of characteristic 0, and let $X$ be an abelian variety over $L$ of dimension $d$. Let $F\in\operatorname{D}^{\operatorname{b}}_{\operatorname{coh}}(X)$ be an object. Then, the following conditions are equivalent:
  1. (1)
@@
  2. (2)
    There is a sub-abelian variety $i:Y\hookrightarrow X$, a rational point $x\in X(L)$, a simple semi-homogeneous vector bundle $V$ on $Y$, and an integer $r$ such that $F\simeq(t_{x})_{*}i_{*}V[r]$, where $t_{x}:X\longrightarrow X$ is the translation map by $x$.
In our variant of this result, the condition $\dim{\rm Ext}^{1}(F,F)\leq d$ will be replaced by a condition on the dimension of a certain group scheme, which we will construct now. Let $L$ be an algebraically closed field of any characteristic, and let $X$ be an abelian variety over $L$ of dimension $d$. Let $s\mathcal{D}$ be the moduli stack of simple relatively perfect universally gluable complexes for the morphism $X\longrightarrow\operatorname{Spec}L$. This is an algebraic stack quasi-separated and locally of finite type over $L$ ([Lie06], [Sta, Tag 0DPX]), and it’s a $\mathbb{G}_{m}$}-gerbe over the coarse space $sD$ which is an algebraic space quasi-separated and locally of finite type over $L$ ([LMB00, Corollaire 10.8]). Given a simple gluable complex $F\in\operatorname{D}^{\operatorname{b}}_{\operatorname{coh}}(X)$, it defines an $L$-point $F]:\operatorname{Spec}L\longrightarrow s\mathcal{D}$, or an $L$-point $F]:\operatorname{Spec}L\longrightarrow sD$. Similarly to [dJO22, 5.8], we consider the action map of $X\times_{L}\widehat{X}$ on $sD$ denoted as $\mu_{F}:X\times_{L}\widehat{X}\longrightarrow sD$ which acts on the $L$-points as $(x,L)\mapsto(t_{x}^{*}F)\otimes L$. We define the stabilizer $S_{F}$ by the cartesian square:
@@
By the following lemma, $S_{F}$ is a closed subgroup variety of $X\times_{L}\widehat{X}$:
@@
With this preparation, we can now state our variant of Proposition 4.18 in all characteristics:
@@
Proof.
This proposition can be proved by a minor modification of the proof of one implication of Proposition 4.18 in [dJO22]. As mentioned in [dJO22, 1.3], the characteristic 0 assumption in this proof can be dropped except for the section [dJO22, 6.1]. In this step, de Jong and Olsson use the condition ${\rm Ext}^{1}(F,F)\leq d$ as well as the characteristic assumption to deduce that $\dim S_{F}\geq d$, and this is the only place where the assumption ${\rm Ext}^{1}(F,F)\leq d$ is used, so the rest of the proof goes through if we instead assume $\dim S_{F}\geq d$. ∎
Remark 4.21.
Let us briefly compare the conditions $\dim S_{F}\geq d$ and ${\rm Ext}^{1}(F,F)\leq d$ for an object $F\in\operatorname{D}^{\operatorname{b}}_{\operatorname{coh}}(X)$ such that ${\rm Ext}^{<0}(F,F)=0$ and ${\rm Ext}^{0}(F,F)=L$. First, we notice that the former condition is stronger than the latter by the combination of Proposition 4.20 and the implication $(2)\implies(1)$ of Proposition 4.18 (which holds in all characteristics). As discussed in [dJO22, 6.1], the tangent space of $S_{F}$ at the origin is the kernel of a certain map $\kappa_{F}:HH^{1}(X)\to{\rm Ext}^{1}(F,F)$. Since the dimension of $HH^{1}(X)$ is $2d$, the condition ${\rm Ext}^{1}(F,F)\leq d$ implies that the dimension of the tangent space is at least $d.$ In characteristic 0, the group variety $S_{F}$ is reduced and hence smooth, so this would imply $\dim S_{F}\geq d$, making the two assumptions equivalent.
@@
[…]
@@
The proposition implies in particular that if $A$ and $B$ are derived equivalent ordinary abelian varieties over $k$, then $A_{\operatorname{can},\overline{\eta}},B_{\operatorname{can},\overline{\eta}}$ are also derived equivalent.
@@
When $\mathcal{X}=A_{\operatorname{can}}$ for an ordinary abelian variety $A$, the condition in the conjecture regarding the Hodge filtration is vacuous. To see this, by the compatibility of Chern classes with cycle classes ([Gro58, (16)]), it suffices to check that the crystalline Chern classes always land in the middle Hodge filtration under the identification $H_{\rm cris}^{*}(A/W)\simeq H_{\rm dR}^{*}(A_{\operatorname{can}}/W)$. By the proof of Proposition 4.16, the isomorphism identifies the Hodge filtration with the slope filtration, but the $i$-th crystalline Chern class always lands in the $i$-th filtered piece of the slope filtration ([LO15, §2.2]), so we have verified the condition. Lemma 4.22 yields the following special case of this conjecture (which in particular applies to kernels of Fourier-Mukai equivalences by the proof of Proposition 4.23):
@@
${\bigoplus_{i=0}^{g-1}H_{\rm dR}^{2i+1}(\mathcal{X}_{\eta}/K)(i)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$$\scriptstyle{\Phi^{\rm dR}_{\mathcal{E}_{\eta}}}\scriptstyle{\simeq}$${\bigoplus_{i=0}^{g-1}H_{\rm dR}^{2i+1}(\mathcal{Y}_{\eta}/K)(i)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\bigoplus_{i=0}^{g-1}H_{\rm cris}^{2i+1}(\mathcal{X}_{s}/K)(i)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi^{\rm cris}_{\mathcal{E}_{s}}}\scriptstyle{\simeq}$${\bigoplus_{i=0}^{g-1}H_{\rm cris}^{2i+1}(\mathcal{Y}_{s}/K)(i)}$